ABC is an equilateral triangle in which BC=2cm find perpendicular distance from A to BC
Answer by Quadratic1600(28)

Hi
So we know that all sides are the same so AB - BC - AC are all 2cm
If you draw a line from A perpendicularly to BC it will meet at a 90 degree angle. So lets call the point where the line from A meets BC (E)
So BE = 1 and CE = 1
So we have the hypotenuse of either AC or AB depending which side of the triangle you choose to work on - It doesn't matter really as they are both 2
So a^2 + b^2 = c^2 which is also the adjacent^2 + opposite^2 = Hypotenuse^2
We know the hypotenuse so lets rearrange so we get
(2^2) (which is the hypotenuse) - (1^2) (which is the line BE or CE depending on which side you are drawing it) = AE
Ok so
4 - 1 = 3
Now SQR of 3 = AE
You will get a positive and negative value as you are square rooting however you can't have a minus side so it will be the positive value.
Hope this helps